a clock shows 12 noon on this monday..
if it loses 1/2% time in the first week and 1/4% time in the next week.
then what will be time after 2 weeks on monday

Answer by edjones(8007)

7*24*.005=.84 hrs
7*24*.0025=.24 hrs
.84+.24=1.08 hrs
60*.08=4.8 min
.8*60=48 sec
12:00:00
-1:04:48
---------
10:55:12
.
Ed